Status Feed Status Information. en-us Copyright 2009 WhiteLabel 1 Log Entry 1000033167 <p>Mon, 19 Jun 2017 19:24 GMT - We are currently experiencing an issue with our VPS platform which is showing some servers as unavailable. More information to follow.</p><p>Mon, 19 Jun 2017 20:52 GMT - Our engineers are currently investigating this issue further.</p><p>Mon, 19 Jun 2017 21:21 GMT - We are currently experiencing an issue with the storage platform on two VS Clusters, our engineers are currently working to resolve this issue however there is no ETTR at this time.</p><p>Mon, 19 Jun 2017 21:56 GMT - Storage drive reconstruction is ongoing, next update will be once reconstruction completes.</p><p>Tue, 20 Jun 2017 01:17 GMT - Drive reconstruction has completed but this has not solved the storage issue. Our engineers are currently working with our third party storage provider to resume services as quickly as possible. No ETTR is available at this time.</p><p>Tue, 20 Jun 2017 04:22 GMT - Following an investigation by our storage partner, we are running a proprietary disk repair against the storage aggregate, it is too early to advise as to a completion time at this stage but we will provide updates during the course of the day. Please accept our apologies for this interruption to service.</p><p>Tue, 20 Jun 2017 07:57 GMT - The disk repair process is ongoing and at this stage we do not have an estimated time for this to complete. Further updates will be provided when information becomes available.</p><p>Tue, 20 Jun 2017 08:48 GMT - The disk repair process is expected to take some time to complete. If you are affected by this issue please contact our Server team on 0333 0142 700 to discuss your options.</p><p>Tue, 20 Jun 2017 11:18 GMT - Our Systems teams are investigating an alternative approach to bring the affected servers online whilst the disk repair is taking place. In the meantime please contact our Server team on 0333 0142 700 to discuss your options.</p><p>Tue, 20 Jun 2017 12:27 GMT - Our engineers are working to bring the backup storage arrays whilst the disk repair continues on the primary array. This process will bring the affected Virtual Servers online with data from the backup taken on the 18th June. We expect this process to complete within a few hours and there will be a performance drop until we can bring the main raid array back online and switch your primary data source back to the optimised platform.</p><p>Tue, 20 Jun 2017 15:17 GMT - In addition to restoring our primary storage aggregate and bringing VPS servers online on the backup storage, we have also started migrating VPS servers onto a newer cluster. This will allow us to balance resources across all our platforms and provide a more reliable service.</p><p>Tue, 20 Jun 2017 17:53 GMT - Our engineers are still working on the migrations to the new cluster and also the primary storage restorations. This work will continue throughout the night however there is no timeframe regarding this. Please accept our apologies for this interruption to service. Further updates will be provided when information becomes available.</p><p>Tue, 20 Jun 2017 20:41 GMT - The migration and restoration work is ongoing and will be throughout the night and onwards. Automatic emails are being sent out to the account related to each server once the migration is complete. Unfortunately there is no ETA on this due to the varied size of each server being relative to its restore time.</p><p>Wed, 21 Jun 2017 06:43 GMT - The migration and restoration work has been ongoing throughout the night and is continuing this morning. A update regarding the current status has been requested from the engineers and we will update this post once more information is known.</p><p>Wed, 21 Jun 2017 08:01 GMT - Server migrations have progressed with a number of servers now online. The process to restore on our primary storage aggregate continues. Further updates will be provided when information becomes available.</p><p>Wed, 21 Jun 2017 10:48 GMT - We are currently in the process of migrating servers onto a newer cluster. The length of time these migrations take are mostly dependant on the amount of data stored on the server and we are doing all we can to restore service to those affected as quickly as possible.</p><p>Wed, 21 Jun 2017 15:26 GMT - The server migrations have continued throughout the day and we have been working to increase the speed of this process wherever possible. This work will continue throughout the night with automatic emails being sent out to the account related to each server once the migration is complete.</p><p>Wed, 21 Jun 2017 19:01 GMT - Migrations have been queued to run overnight and will continue through the morning. Automatic emails are being sent out to the account related to each server once the migration is complete.</p><p>Thu, 22 Jun 2017 10:17 GMT - The proprietary disk repair process on our primary storage aggregate is ongoing and will take a considerable time to complete. Server migrations are continuing and we have been working to increase the efficiency of this process.</p><p>Thu, 22 Jun 2017 13:36 GMT - Migrations onto the new cluster are ongoing and we will contact customers as servers are brought back online.</p><p>Fri, 23 Jun 2017 07:28 GMT - Our work on speeding up the migration process has been successful, migrations of servers onto the new cluster are progressing well and a significant proportion of servers are now online. We will continue to contact customers as servers are brought back online.</p><p>Fri, 23 Jun 2017 11:23 GMT - The majority of customers are now back online with their servers migrated onto an alternative cluster. The migration of the remaining servers is ongoing.</p><p>Fri, 23 Jun 2017 14:59 GMT - The migration of servers onto an alternative cluster continues, over 50% have are now online. The length of time for migrations to complete is dependent on the size of the storage drives for each server.</p><p>Sat, 24 Jun 2017 13:50 GMT - Migrations are continuing. The majority of servers have been migrated and are now online with the remaining either being queued or needing manual amendments to complete. Users are receiving notification emails once their server has been migrated. The length of time for migrations to complete is dependent on the size of the storage drives for each server.</p><p>Sun, 25 Jun 2017 11:13 GMT - The majority of the affected VPS servers have now successfully migrated across and are online. Users are receiving notification emails once their server has been migrated. Work is ongoing to migrate the remaining servers. The length of time for migrations to complete is dependent on the size of the storage drives for each server.</p><p>Mon, 26 Jun 2017 06:22 GMT - The vast majority of servers are back online and we are progressing with the remaining migrations If your server is not accessible and you have not yet been in contact with us, please call us on 0333 0142 700 so we can look at your individual server and discuss possible options.</p><p>Mon, 26 Jun 2017 10:18 GMT - This issue is resolved. We are aware of a small number of fringe cases where servers may be offline or poorly performing, and we are in contact with the individual customers affected. If your VPS is still experiencing an issue and you have not heard from us recently, please call us on 0333 0142 700.</p> Mon, 19 Jun 2017 19:24:56 GMT /StatusLogEntry.php?StatusLogEntry_id=StatusLogEntry%3A1000033167